# Enhancing Data Lakes with GraphAr: Efficient Graph Data Management with a Specialized Storage Scheme

### v3

@@ 5 Optimized Label Filtering
5.1 Handling Simple Conditions
@@
Definition 3 (Simple Condition Filtering)
Given a label $l$, and an existence/absence indicator $e$, the simple label filtering returns the PAC $\mathcal{C}$, where
$v\in\mathcal{C},\text{ if }v.label[l]=e\\ v\notin\mathcal{C},\text{ if }v.label[l]\neq e\end{aligned}\right.$ (1)
@@
5.2 Extending to Complex Conditions
@@
Definition 4 (Complex Condition Filtering)
@@
$v\in\mathcal{C},\text{ if }f(v)=\text{true}\\ v\notin\mathcal{C},\text{ if }f(v)=\text{false}\end{aligned}\right.$ (2)
The intuitive approach would be to tackle each vertex independently, decoding the RLE format into the binary representation detailed earlier. However, a direct evaluation of the UDF for every vertex proves impractical, as it retains the same complexity as the most straightforward approach.
@@
$\forall u,v\in[s,e),l\in[0,k)\quad u.label[l]=v.label[l].$ (4)
Consequently, it is sufficient to call the UDF for the vertex $s$ alone, as for any vertex $v$ in the interval $[s,e)$, $v$ and $s$ share the same labels thus $f(v)=f(s)$.
Figure 7: An example of merging intervals.
Merge-based algorithm. Partitioning an interval into multiple segments proves unnecessary and counterproductive as it would escalate complexity. Therefore, our focus narrows down to intervals formed by existing positions, which also addresses the second question. To obtain the exact intervals, we can sort the positions in all interval lists $P_{0},P_{1},\dots,P_{k-1}$. This sorting can be accelerated by leveraging the inherent order within the $k$ lists, allowing for seamless merging of $k$ sorted lists into one list $P$. Figure 7 demonstrates an example of interval determination for a complex condition containing two labels. Within the interval $[P[i],P[i+1])$, the vertices share identical labels, necessitating the UDF to be invoked solely for one representative vertex. Additionally, the presence of position $P[6]$ for both labels underscores the importance of merging to avoid redundant computations.
@@
6 Evaluation
In this section, we evaluate GraphAr on a range of real-world and synthetic graphs, through micro-benchmarks focusing on neighbor retrieval and label filtering, as well as end-to-end workloads. Our findings validate GraphAr as an efficient storage scheme for LPG storage and querying in data lakes.
Table 1: Statistics of the graphs in our evaluation.
Abbr. Graph $\mathbf{|V|}$ $\mathbf{|E|}$
AR arabic-2005 [29] 22.7M 1.27B
BL bloom [14] 33.0K 29.7K
CI citations [14] 264K 221K
CL cont1-l [56] 1.92M 7.03M
DM degme [56] 659K 8.13M
HW hollywood-2009 [29] 1.14M 113M
OL icij-offshoreleaks [14] 1.97M 3.27M
PP icij-paradise-papers [14] 163K 364K
IC indochina-2004 [29] 7.41M 384M
LR LargeRegFile [56] 2.11M 4.94M
NM network-management [14] 83.8K 181K
AX ogbn-arxiv [39] 169K 1.17M
MA ogbn-mag [39] 736K 21.1M
OS openstreetmap [14] 71.6K 76.0K
OR orkut [44] 3.07M 213M
PO pole [14] 61.5K 105.8K
SF30 SNB Interactive SF-30 [32] 99.4M 655M
SF100 SNB Interactive SF-100 [32] 318M 2.15B
SF300 SNB Interactive SF-300 [32] 908M 6.29B
TP tp-6 [56] 1.01M 10.7M
TT twitter-trolls [14] 281K 493K
U2 uk-2002 [29] 18.5M 589M
U5 uk-2005 [29] 39.5M 1.85B
WB webbase-2001 [29] 118M 2.01B
WK wiki [42] 13.6M 437M
6.1 Experimental Setup
Platform. Experiments are conducted on an Alibaba Cloud r6.6xlarge instance, equipped with a 24-core Intel(R) Xeon(R) Platinum 8269CY CPU at 2.50GHz and 192GB RAM, running 64-bit Ubuntu 20.04 LTS. The data is hosted on a 200GB PL0 ESSD with a peak I/O throughput of 180MB/s. Additional tests on other platforms and S3-like storage yield similar results. For timing metrics, we use single-threaded executions and report either average or distribution times based on multiple runs for accuracy.
Baselines. GraphAr is developed in C++ on Apache Arrow [2], an open-source, high-performance library that supports columnar formats like Parquet and ORC. We benchmark GraphAr against Apache Arrow/Parquet (version 13.0.0), due to the popularity and high-performance of Parquet. While there are other related works in this area, we discuss them in Section 7 and do not directly compare them with GraphAr, since they are either not applicable or demonstrate worse performance than Parquet in the context of data lakes. Unless otherwise specified, both GraphAr and the baseline follow Parquet’s default configurations, which include a row group length of $1024\times 1024$ rows and a 1MB page size.
Datasets. As summarized in Table 1, our evaluation includes a variety of graphs, span different sizes and domains, including social networks and web graphs. We also use synthetic graphs generated by the LDBC SNB data generator [32], designed to mimic real-world graph characteristics. They are part of the LDBC Social Network Benchmark [32, 57], and we use its query set for end-to-end workload assessments.
6.2 Micro-Benchmark of Neighbor Retrieval
6.2.1 Storage Efficiency
We compare GraphAr with baseline methods by measuring the storage consumed by encoded Parquet files that store the graph’s topological data. Two baseline approaches are considered: 1) “plain”, which employs plain encoding for the source and destination columns, and 2) “plain + offset”, which extends the “plain” method by sorting edges and adding an offset column to mark each vertex’s starting edge position. As Figure 7(a) depicts, the inclusion of offsets results in a modest increase in storage requirements, with an increase in space usage ranging from $0.5\%$ to $14.8\%$, as the number of vertices is typically much smaller than the number of edges.
In contrast, GraphAr leverages delta encoding for source and destination columns and utilizes plain encoding for the offsets, adhering to default Arrow/Parquet settings. The result is a notable storage advantage: on average, GraphAr requires only 27.3% of the storage needed by the baseline “plain + offset”. This efficiency in storage is particularly beneficial for query performance, as data lake queries are often I/O-bound. The transition from storage efficiency to retrieval performance is elaborated further in the next experiment.
@@ 6.2.1 Storage Efficiency
(c) Data transformation time.
Figure 8: Micro-benchmark of neighbor retrieval.
6.2.2 Performance of Neighbor Retrieval
@@
To evaluate GraphAr’s efficiency in neighbor retrieval, we query vertices with the largest degree in selected graphs, maintaining edges in CSR-like or CSC-like formats depending on the degree type. Figure 7(b) shows that GraphAr significantly outperforms the baselines, achieving speedups ranging from 49.8×49.8$\times$49.8 × to 9232.4×9232.4$\times$9232.4 × over the “plain” method and 2.1×2.1$\times$2.1 × to 6.1×6.1$\times$6.1 × over the “plain + offset” baseline. These gains are attributed to the offset integration and delta encoding, as well as our proposed BMI-based decoding strategy. The offset integration alone accounts for a speedup of between 11.0×11.0$\times$11.0 × and 4177.2×4177.2$\times$4177.2 ×, and delta encoding contributes an additional 1.8×1.8$\times$1.8 × to 3.6×3.6$\times$3.6 × speedup on top of that. While not explicitly shown in the log-scaled Figure 7(b), the use of BMI and SIMD further enhances performance, with an average improvement of $20.1\%$.
6.2.3 Performance of Data Transformation
Given that GraphAr is designed for storing LPGs in data lakes, the efficiency of converting original graph data into the GraphAr format is crucial. We focus on the time required for this transformation, using four large graphs (U2, AR, WB and U5) represented as Arrow Tables, a standardized in-memory format in big data systems. Since graphs generally have significantly more edges than vertices and GraphAr employs CSR/CSC-like layouts requiring edge sorting, generating topological data becomes the most time-intensive part.
To dissect this time overhead, we perform a breakdown analysis in Figure 7(c), which shows that we can generate topological data for over $3.9$ million edges per second. The process involves three steps: 1) sorting the edges using Arrow/Acero’s internal order_by operator, labeled as “sort”; 2) generating vertex offsets, labeled as “offset”; and 3) writing the sorted and offset data into Parquet files with specific encoding, labeled as “output”. Generally, this process has a time complexity of $O(|E|log|E|)$ sequentially. Considering this transformation is a one-time, offline operation that substantially reduces future data retrieval times, the associated overhead is acceptable. The most time-consuming step, sorting, offers room for further optimization through distributed frameworks like Spark.
6.3 Micro-Benchmark of Label Filtering
This section evaluates GraphAr’s efficiency in storing and filtering vertex labels. We employ datasets from OGB [39] and Neo4j [14], which feature property graphs with multiple vertex labels, with the number of labels (ranging from $3$ to $349$) indicated under the graph name in Figure 8(a).
6.3.1 Storage Efficiency
We assess storage efficiency by measuring the size of encoded Parquet files used for storing vertex labels. Two baseline methods serve for comparison: The first, termed “string,” concatenates all labels of a vertex into a single string column using BYTE_ARRAY datatype and plain encoding. The second, named “binary (plain),” represents each label in a separate binary column using BOOLEAN datatype and plain encoding. Our approach, denoted as “binary (RLE),” further optimizes this by utilizing Run-Length Encoding (RLE).
As shown in Figure 8(a), our RLE-based method substantially outperforms the baselines, requiring on average only $2.9\%$ and $10.1\%$ of the storage space compared to the “string” and “binary (plain)” methods, respectively. While Parquet does support dictionary encoding that could potentially enhance the “string” baseline, we excluded this from our evaluation. The reason being that, despite some storage gains, dictionary encoding incurs a decoding slowdown of up to 10×10$\times$10 × due to the extra overhead of storing the dictionary, especially when label numbers are high. Our RLE-based strategy strikes a balance between storage efficiency and decoding performance, as demonstrated in the subsequent experiments.
6.3.2 Performance of Simple Condition Filtering
Recognizing that filtering based on simple conditions represents the cornerstone operation in graph query languages, we prioritize evaluating this operation. For each graph, we perform experiments where we consider each label individually as the target label for filtering, and determine the vertices with that label. To ensure accuracy, each experiment is repeated for $100$ times and the total execution time is reported.
Figure 8(b) illustrates the results, demonstrating that the GraphAr method significantly improves the performance of label filtering based on simple conditions. The most straightforward approach, “string”, which involves decoding the string of labels and conducting matching for each vertex, is the slowest. The “binary (plain)” method separates labels into individual columns and utilizes a binary representation, while the ‘binary (RLE)” method further optimizes the encoding by using RLE. However, both of these methods still require evaluating each vertex. In contrast, our method, referred to as “binary (RLE) + interval”, simply selects all satisfied intervals.
In Figure 8(b), for each graph, we report the middle value of the execution time among filtering each label as the height of the bar, with the error bar representing the range of execution time. Our method may have a large range on some graphs (AX, MA) due to the varying encoding efficiency (i.e., the number of intervals generated) for different columns. However, since the number of intervals is not larger than the number of vertices in any case, our method consistently outperforms the baselines. On average, it achieves a speedup of 6.0×6.0$\times$6.0 × over the “string” method, $\times3$ over the “binary (plain)” method, and 2.3×2.3$\times$2.3 × over the “binary (RLE)” method.
@@ 6.3.2 Performance of Simple Condition Filtering
Figure 9: Micro-benchmark of label filtering.
6.3.3 Performance of Complex Condition Filtering
@@
We also assess the performance of label filtering based on complex conditions using the same graphs as mentioned above. For graphs obtained from Neo4j, we first refer to the provided documentation to identify a filtering operation that involves two labels. If not provided, we create a condition by combining two related labels using either the logical AND operator (if there are vertices satisfying the condition) or OR (otherwise) to reflect real-world semantics. For graphs obtained from OGB, we utilize the first two labels and combine them using OR as the filtering condition.
@@
Figure 8(c) presents the speedup of different methods against the most straightforward approach (referred to as “string” in Figure 8(a) and Figure 8(b)). The results demonstrate that GraphAr performs the best for all test cases. Further analysis reveals that the performance improvement is attributed to the binary representation (as observed in the comparison between “binary (plain)” and “string”), and utilization of RLE (as seen in the comparison between “binary (RLE)” and “binary (plain)”). However, the most significant improvement comes from the merge-based decoding, as observed in the comparison between “binary (RLE) + merge” and “binary (RLE)”, which results in a further speedup of up to 56.3×56.3$\times$56.3 ×.
6.3.4 Scale Up the Number of Labels
@@
Figure 10 illustrates the execution time of filtering conditions with varying numbers of labels, focusing on BL and MA, which are selected from different datasets (Neo4j and OGB) and have a relatively large number of labels. We test the filtering with $i$ labels by combining the first $i$ labels in a graph using the OR operator as the condition, to ensure that there are vertices satisfying the condition. To reduce experimental noise, we conduct the experiments for $100$ times and report the total execution time for BL and the average time for MA.
@@
As shown in the figure, GraphAr consistently outperforms others on BL. While on MA, it performs best when the number of involved labels is no more than $40$. As the number of labels continuously increases, it performs worse than the baseline “binary (RLE)”, which is due to the number of merged intervals also increasing. In the worst case, the UDF is called for each vertex, means any two consective vertices have different labels. Considering the overhead of merging intervals from different columns, our method may perform worse than directly evaluating the UDF for each vertex. Fortunately, our investigation of real-world workloads reveals that the number of filtered labels in user-written queries is often limited. For instance, in the documentation of Neo4j examples  [14], the filtering involves no more than $5$ labels. Consequently, the performance of our method remains highly promising.
Figure 10: Execution time with the number of labels increases on, tested on BL (left) and MA (right).
6.4 Storage Media

### v4

@@ 5. Optimized Label Filtering
5.1. Handling Simple Conditions
@@
Definition 0 (Simple Condition Filtering).
Given a label $l$ and an existence/absence indicator $e$, the simple condition label filtering returns the PAC $\mathcal{C}$, where
(1) $v\in\mathcal{C},\text{ if }v.label[l]=e\\ v\notin\mathcal{C},\text{ if }v.label[l]\neq e\end{aligned}\right.$
@@
5.2. Extending to Complex Conditions
@@
Definition 0 (Complex Condition Filtering).
@@
(2) $v\in\mathcal{C},\text{ if }f(v)=\text{true}\\ v\notin\mathcal{C},\text{ if }f(v)=\text{false}\end{aligned}\right.$
The intuitive approach would be to tackle each vertex independently, decoding RLE into the binary representation. However, directly evaluating the UDF for every vertex proves impractical, as it retains the same complexity as the most straightforward approach.
@@
(4) $\forall u,v\in[s,e),l\in[0,k)\quad u.label[l]=v.label[l].$
Consequently, it is sufficient to call the UDF for the vertex $s$ alone, as for any vertex $v$ in the interval $[s,e)$, $v$ and $s$ share the same labels, thus $f(v)=f(s)$.
Figure 7. An example of merging intervals.
Merge-based algorithm. Partitioning an interval into multiple segments proves unnecessary and counterproductive as it would escalate complexity. Therefore, our focus narrows down to intervals formed by existing positions, which also addresses the second question. To obtain the exact intervals, we can sort the positions in all interval lists $P_{0},P_{1},\dots,P_{k-1}$. This sorting can be accelerated by leveraging the inherent order within the $k$ lists, allowing for seamless merging of $k$ sorted lists into one list $P$. Figure 7 demonstrates an example of interval determination for a complex condition containing two labels, Asian and Enrollee. Within the interval $[P[i],P[i+1])$, the vertices share identical labels, necessitating the UDF to be invoked solely for one representative vertex. Additionally, the presence of position $P[6]$ for both labels underscores the importance of merging to avoid redundant computations.
@@
6. Evaluation
In this section, we evaluate GraphAr on a range of graphs, through micro-benchmarks and end-to-end graph query workloads. We also illustrate the potential benefits GraphAr offers to current graph processing systems. Our findings validate GraphAr as an efficient storage scheme for LPG storage and querying in data lakes.
@@ 6. Evaluation
(c) Transformation time.
Figure 8. Micro-benchmark of neighbor retrieval.
Table 1. Statistics of the graphs in our evaluation.
Abbr. Graph $\mathbf{|V|}$ $\mathbf{|E|}$
A5 Alibaba synthetic (scale 5) 75.0M 4.93B
A7 Alibaba synthetic (scale 7) 100M 6.69B
AR arabic-2005 (Davis and Hu, 2011) 22.7M 1.27B
BL bloom (neo, 2024c) 33.0K 29.7K
CF com-friendster (Leskovec and Krevl, 2014) 65.6M 1.81B
CI citations (neo, 2024c) 264K 221K
CL cont1-l (Rossi and Ahmed, 2015) 1.92M 7.03M
DM degme (Rossi and Ahmed, 2015) 659K 8.13M
G8 Graph500-28 (gra, 2024) 268M 4.29B
G9 Graph500-29 (gra, 2024) 537M 8.59B
HW hollywood-2009 (Davis and Hu, 2011) 1.14M 113M
OL icij-offshoreleaks (neo, 2024c) 1.97M 3.27M
PP icij-paradise-papers (neo, 2024c) 163K 364K
IC indochina-2004 (Davis and Hu, 2011) 7.41M 384M
NM network-management (neo, 2024c) 83.8K 181K
AX ogbn-arxiv (Hu et al., 2020) 169K 1.17M
MA ogbn-mag (Hu et al., 2020) 736K 21.1M
OS openstreetmap (neo, 2024c) 71.6K 76.0K
PO pole (neo, 2024c) 61.5K 105.8K
SF30 SNB Interactive SF-30 (Erling et al., 2015) 99.4M 655M
SF100 SNB Interactive SF-100 (Erling et al., 2015) 318M 2.15B
SF300 SNB Interactive SF-300 (Erling et al., 2015) 908M 6.29B
TP tp-6 (Rossi and Ahmed, 2015) 1.01M 10.7M
TT twitter-trolls (neo, 2024c) 281K 493K
U2 uk-2002 (Davis and Hu, 2011) 18.5M 589M
U5 uk-2005 (Davis and Hu, 2011) 39.5M 1.85B
WB webbase-2001 (Davis and Hu, 2011) 118M 2.01B
WK wiki (Kunegis, 2013) 13.6M 437M
6.1. Experimental Setup
Platform. If not otherwise mentioned, our experiments are conducted on an Alibaba Cloud r6.6xlarge instance, equipped with a 24-core Intel(R) Xeon(R) Platinum 8269CY CPU at 2.50GHz and 192GB RAM, running 64-bit Ubuntu 20.04 LTS. The data is hosted on a 200GB PL0 ESSD with a peak I/O throughput of 180MB/s. Additional tests on other platforms and S3-like storage yield similar results. For timing metrics, we use single-threaded executions and report either average or distribution times based on multiple runs for accuracy. Exceptionally, the integration experiments utilize a cluster of 8 separate instances to emulate a distributed environment.
Baselines. GraphAr is developed in C++ on Apache Arrow (arr, 2024), an open-source, high-performance library that supports columnar formats like Parquet and ORC. For the micro-benchmarks, we compare GraphAr against Arrow/Parquet (version 13.0.0), due to the popularity and high-performance of Parquet. Both GraphAr and the baseline follow Parquet’s default configurations, which include a row group length of $1024\times 1024$ and a 1MB page size. For end-to-end workloads, we compare GraphAr against widely-used frameworks including Apache Acero (Ace, 2024), Apache Pinot (pin, 2024) and Neo4j (Community 5.21.0) (neo, 2024b). In the integration experiments, we incorporate GraphAr into GraphScope (Fan et al., 2021), a widely-used graph processing system, and compare its integrated performance against GraphScope’s original implementation.
Datasets. As summarized in Table 1, our evaluation includes a variety of graphs, span different sizes and domains, including social networks and web graphs. We also use synthetic graphs generated by data generators of the LDBC SNB (Erling et al., 2015) and Graph500 (gra, 2024), both of which are widely recognized benchmarks. Additionally, we utilize graphs (A5 and A7) generated that closely mimic the characteristics of graphs in the e-commerce production environment at Alibaba.
6.2. Micro-Benchmark of Neighbor Retrieval
Storage efficiency. We compare GraphAr with baseline methods by measuring the storage consumed by encoded Parquet files that store the graph’s topological data. Two baseline approaches are considered: 1) “plain”, which employs plain encoding for the source and destination columns, and 2) “plain + offset”, which extends the “plain” method by sorting edges and adding an offset column to mark each vertex’s starting edge position. As Figure 8(a) depicts, the inclusion of offsets results in a modest increase in storage requirements, with space usage growing by $0.5\%$ to $14.8\%$, as the number of vertices is typically much smaller than the number of edges.
GraphAr leverages delta encoding for source and destination columns and plain encoding for offsets. The result is a notable storage advantage: on average, it requires only 29.2% of the storage needed by the baseline “plain + offset”. This efficiency in storage is particularly beneficial for query performance, as data lake queries are often I/O-bound. The transition from storage efficiency to retrieval performance is elaborated further in the next experiment.
@@ 6.2. Micro-Benchmark of Neighbor Retrieval
Performance of neighbor retrieval. To evaluate GraphAr’s efficiency in neighbor retrieval, we query vertices with the largest degree in selected graphs, maintaining edges in CSR-like or CSC-like formats depending on the degree type. Figure 8(b) shows that GraphAr significantly outperforms the baselines, achieving an average speedup of 4452×4452$\times$4452 × over the “plain” method, 3.05×3.05$\times$3.05 × over “plain + offset”, and 1.23×1.23$\times$1.23 × over “delta + offset”. These gains are attributed to the offset integration and delta encoding, as well as our BMI-based decoding. The offset integration alone accounts for an average speedup of 1993×1993$\times$1993 ×, and delta encoding provides an additional 2.48×2.48$\times$2.48 × speedup. Our innovative decoding method, which leverages BMI and SIMD, further enhances performance within this optimized context, achieving a 1.23×1.23$\times$1.23 × speedup on top of “delta + offset”.
@@
(d) Scaling the number of labels, tested on BL (left) and MA (right).
Figure 9. Micro-benchmark of label filtering.
Performance of data transformation. Given that GraphAr is designed for storing LPGs in data lakes, the efficiency of converting original graph data into the GraphAr format is crucial. Graphs generally have significantly more edges than vertices and GraphAr employs CSR/CSC-like layouts requiring edge sorting. Thus, generating topological data becomes the most time-intensive part. To assess the overhead, we analyze the time to convert four real-world graphs (U2, AR, WB, and U5), each initially in the form of Arrow Tables, a standardized in-memory format in big data systems.
Figure 8(c) illustrates the time breakdown. The process involves three steps: 1) sorting the edges, using Arrow’s order_by operator, labeled as “sort (Arrow)”; 2) generating vertex offsets, labeled as “offset”; and 3) writing the sorted and offset data into Parquet files with specific encoding, labeled as “write”. The sorting step is most time-consuming, which has an average time complexity of $O(|E|log|E|)$ when executed sequentially, while the steps of generating offset and writing with encoding both have a lower time complexity of $O(|E|)$. For further optimization, we leverage the parallel sorting algorithms provided by Intel(R) Threading Building Blocks (TBB, 2024), which significantly reduces the sorting time, labeled as “sort (TBB)”. By employing 24 threads in our test setup, the sorting time is only $8.9\%$ of the original, on average. Given this transformation is a one-time, offline operation that substantially reduces future data retrieval times, the associated overhead—which is within 1 minute for generating topological data for over 1 billion edges—is acceptable.
6.3. Micro-Benchmark of Label Filtering
This section evaluates GraphAr’s efficiency in storing and filtering vertex labels. We employ datasets from OGB (Hu et al., 2020), Neo4j (neo, 2024c) and Alibaba’s synthetic data generator, which feature property graphs with multiple vertex labels, with the number of labels (ranging from $3$ to $349$) indicated under the graph name in Figure 9(a).
Storage efficiency. We assess storage efficiency by measuring the size of encoded Parquet files used for storing vertex labels. Two baseline methods serve for comparison: The first, termed “string”, concatenates all labels of a vertex into a single string column using BYTE_ARRAY datatype and plain encoding. The second, named “binary (plain)”, represents each label in a separate binary column using BOOLEAN datatype and plain encoding. Our approach, denoted as “binary (RLE)”, further optimizes this by utilizing RLE.
As shown in Figure 9(a), our RLE-based method substantially outperforms the baselines, requiring on average only $2.5\%$ and $8.4\%$ of the storage space compared to the “string” and “binary (plain)” methods, respectively. While Parquet does support dictionary encoding that could potentially enhance the “string” baseline, we excluded this from our evaluation. The reason being that, despite some storage gains, dictionary encoding incurs a decoding slowdown of up to 10×10$\times$10 × due to the extra overhead of storing the dictionary, especially when label numbers are high. Our RLE-based strategy strikes a balance between storage efficiency and decoding performance, as demonstrated in the subsequent experiments.
Performance of simple condition filtering. Recognizing that filtering based on simple conditions represents the cornerstone operation in graph query languages, we prioritize evaluating this operation. For each graph, we perform experiments where we consider each label individually as the target label for filtering, and determine the vertices with that label. For accuracy, each experiment is repeated $100$ times and the total execution time is reported.
Figure 9(b) illustrates the results, demonstrating that GraphAr’s method significantly improves the performance of label filtering based on simple conditions. The most straightforward approach, “string”, which involves decoding the string of labels and conducting matching for each vertex, is the slowest. The “binary (plain)” method separates labels into individual columns and utilizes a binary representation, while the ‘binary (RLE)” method further optimizes the encoding by using RLE. However, both of these methods still require evaluating each vertex. In contrast, our method of “binary (RLE) + interval”, simply selects all satisfied intervals.
In Figure 9(b), for each graph, we report the middle value of the execution time among filtering each label as the height of the bar, with the error bar representing the range of execution time. Our method may have a large range on some graphs (AX, MA) due to the varying encoding efficiency (i.e., the number of intervals generated) for different columns. However, since the number of intervals is not larger than the number of vertices in any case, our method consistently outperforms the baselines. On average, it achieves a speedup of 14.8×14.8$\times$14.8 × over the “string” method, 8.9×8.9$\times$8.9 × over the “binary (plain)” method, and 7.4×7.4$\times$7.4 × over the “binary (RLE)” method.
@@ 6.3. Micro-Benchmark of Label Filtering
Performance of complex condition filtering. We also assess the performance of label filtering based on complex conditions using the same graphs as mentioned above. For graphs obtained from Neo4j, we first refer to the provided documentation to identify a filtering operation that involves two labels. If not provided, we create a condition by combining two related labels using either the logical AND operator (if there are vertices satisfying the condition) or OR (otherwise) to reflect real-world semantics. For other graphs, we combine the first two labels by OR as the filtering condition.
@@
Figure 9(c) presents the performance of different methods, measured as the total execution time of $100$ runs. The results demonstrate that GraphAr performs the best for all test cases. Further analysis reveals that the performance improvement is attributed to the binary representation (as seen in the comparison between “binary (plain)” and “string”), and utilization of RLE (as seen in the comparison between “binary (RLE)” and “binary (plain)”). However, the merge-based decoding method yields the largest gain, where “binary (RLE) + merge” outperforms the “binary (RLE)” method by up to 60.5×60.5$\times$60.5 ×.
@@
Scale up the number of labels. Figure 9(d) illustrates the average execution time of filtering conditions with varying numbers of labels, focusing on BL and MA, which are selected from different datasets (Neo4j and OGB) and have a relatively large number of labels. We test the filtering with $i$ labels by combining the first $i$ labels through OR as the condition.
@@
As shown in the figure, GraphAr consistently outperforms others on BL. While on MA, it performs best when the number of involved labels is no more than $40$. As the number of labels continuously increases, it performs worse than the baseline “binary (RLE)”, which is due to the number of merged intervals also increasing. In the worst case, the UDF is called for each vertex, means any two consecutive vertices have different labels. Considering the overhead of merging intervals from different columns, our method may perform worse than directly evaluating the UDF for each vertex. Fortunately, our investigation of real-world workloads reveals that the number of filtered labels in user-written queries is often limited. For instance, in the Neo4j documentation examples  (neo, 2024c), the filtering involves at most $5$ labels, suggesting our method is highly promising.
6.4. Storage Media